Calculate the acceleration due to gravity given that the radius of earth 'Re is 6.4×106m radius of the moon's orbit is 60.wee and period, T is 27.3 days
1
Expert's answer
2018-04-26T10:21:08-0400
Acceleration due to gravity:
a=(4·π^2·R)/T^2
Where R=60.wee=60·6.4·〖10〗^6=384·〖10〗^6 m
T=27.3 days=27.3·24·60·60=2.36·〖10〗^6 sec
a=(4·〖3.14〗^2·384·〖10〗^6)/〖(2.36·〖10〗^6)〗^2 =0.0027 m/s^2